Find the first five terms of the sequence whose nth term is given by the following. -2n -8
the first term?
the second term?
the third term?
the fourth term?
the fifth term?

Answer by ikleyn(52781)

.
Find the first five terms of the sequence whose nth term is given by the following. -2n -8

the first term?     -2*1 - 8 = -2-8 = -10.        <--- Substirute 1 for n and calculate . . . 
the second term?    -2*2 - 8 = -4-8 = -12         <--- Substitute 2 for n and calculate . . . 
the third term?     continue in the same manner . . . 
the fourth term?
the fifth term?